Friday, 17 June 2011

Update 2pm - 17/6

Engineers have confirmed that routing issues affecting some VDS clients have now been resolved and normal services returned to affected customers.

Work is continuing on the recovery of the remaining shared servers, but we are still unable to provide any accurate indication of time for resolution on these.

Reseller systems are progressing as outlined yesterday. Staff are working with the registries presently to configure and test connections prior to relaunch.

We are aiming to post some more detailed information on the nature of the attack and more detailed status of remaining issues later today, however our focus continues to be on restoration of services and assisting clients as quickly as possible at this stage.